-6r + s - rs r=5 s=7
Answer by RAY100(1637)

-6r +s -rs
.
r(-6-s) +s
.
5(-6-7) +7
.
5(-13) +7
.
-65 +7
.
-58
.
check using direct subst into initial eqn
.
-6(5)+(7) -(5)(7) = -30 +7 -35 = -65 +7 =-58,,,,ok